### v1

# Individual Identification Using Radar-Measured Respiratory and Heartbeat Features

Abstract
This study proposes a method for radar-based identification of individuals using a combination of their respiratory and heartbeat features. In the proposed method, the target individual’s respiratory features are extracted using the modified raised-cosine-waveform model and their heartbeat features are extracted using the mel-frequency cepstral analysis technique. To identify a suitable combination of features and a classifier, we compare the performances of nine methods based on various combinations of three feature vectors with three classifiers. The accuracy of the proposed method in performing individual identification is evaluated using a 79-GHz millimeter-wave radar system with an antenna array in two experimental scenarios and we demonstrate the importance of use of the combination of the respiratory and heartbeat features in achieving accurate identification of individuals. The proposed method achieves accuracy of 96.33% when applied to a five-day dataset of six participants and 99.39% when applied to a public one-day dataset of thirty participants.
Index Terms:
classification, heartbeat features, individual identification, millimeter-wave radar, respiratory features
I Introduction
At present, with the ongoing advancements in Internet of Things (IoT) technology, techniques for sensing of various information types in real environments have been attracting attention. IoT technologies are easing the digitalization of society and enabling the construction of sensing networks for use in daily life [1, 2]. In particular, research into identification of individuals using sensors has drawn attention in a variety of applications. Currently, facial recognition using cameras is primarily being researched and commercialized [3, 4]. Although facial recognition can achieve high-precision identification, it also raises privacy concerns and is unsuitable for sensing applications in private spaces. Additionally, external factors, such as low lighting, can lead to reduced recognition accuracy.
Rather than use cameras, contact sensor-based methods for individual identification that sense fingerprints [5], veins [6], and electrocardiograms (ECGs) [7, 8, 9] have been reported. Furthermore, individual identification techniques using respiratory sounds measured with microphones have also been reported [10, 11]. Han et al. [10] proposed an individual identification method that used respiratory sounds acquired from bone conduction microphones embedded in earphones and achieved identification accuracy of 95.17% for 20 participants. Although these studies focused on use of contact sensors to obtain the physiological signals required for individual identification, it can be difficult to attach the sensors or place the required sensors close to the body, and this factor will also hinder the identification of uncooperative individuals.
Recently, radar has been used widely for measurement of human bodies because both limb movements and physiological signals can be acquired in a noncontact manner and privacy concerns are reduced when compared with use of camera-based sensors [12, 13, 14]. Radar techniques have been used for respiration measurement  [15, 16, 17], heartbeat [18, 19, 21, 20] and also for individual identification applications [22, 23].
Limb motions such as walking and sitting movements are often used to identify individuals, e.g., in individual identification of walking and sitting persons using Wi-Fi signals [24, 25, 26, 27, 28] and radar techniques [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40]. To enable individual identification of a person that is remaining still, Wi-Fi signals and radar have also been used to measure their physiological signals [41, 42, 43, 44, 48, 46, 45, 47, 49]. For example, Rahman et al. [44] extracted features from radar-measured respiration characteristics and identified individuals using a combination of the $k$-nearest neighbor ($k$-NN) algorithm and majority voting. Although respiration measured via radar is useful for individual identification, respiration is also known to have both voluntary and involuntary aspects [50], which raises questions about its effectiveness for use in long-term identification. In contrast, Lin et al. [49] conducted individual identification using heartbeat features that were measured using a pair of radar systems, with accuracy of 98.61% being achieved for 78 participants.
In this paper, we propose an individual identification method that uses both respiratory and heartbeat features for the first time. The contributions of this paper are listed as follows:
  • We propose a method to extract respiratory features in addition to extraction of respiratory intervals, and confirm the effectiveness of using these features to perform individual identification through experiments.
  • We propose a method to extract features of heartbeat harmonic components and demonstrate the effectiveness of use of these features in individual identification.
  • Through a series of experiments conducted over five days on six participants, we evaluate the long-term effectiveness of the proposed method. We also verify the method’s accuracy when using the respiratory and heartbeat features separately and discuss the advantages of combining these features. Additionally, we validate the identification accuracy for larger groups using a public dataset with 30 participants.
II Respiratory features for individual identification
II-A Waveform Feature Extraction Method
If the radar system receives a single dominant echo from a human body, the body’s displacement can be estimated to be
$d(t)=\frac{\lambda}{4\pi}\mathrm{unwrap}\quantity{\angle s(t)},$ (1)
where $s(t)$ is the signal reflected from the target, $\mathrm{unwrap}\{\cdot\}$ is the phase unwrapping operator, and $\angle$ represents the phase of a complex number. When the target person is stationary, $d(t)$ mainly contains the displacement components that result from the person’s respiration and heartbeat.
The respiratory displacement caused by normal breathing can be expressed as a semi-periodic function of time with three phases: the expiratory phase, the expiratory plateau, and the inspiratory phase [51]. Fig. 1 shows an example of a radar-measured displacement waveform that highlights the three phases described above. To extract individual features related to the three respiratory phases, we use the modified raised-cosine-waveform (MRCW) model proposed by Hsieh et al. [52]. In the MRCW model, in addition to the respiratory period and amplitude, the user can also specify the time duration ratio of the expiratory, expiratory plateau, and inspiratory phases.
The MRCW model for one cycle $T=1/f$ can be expressed using six parameters as follows:
$\begin{split}&d_{\mathrm{MRCW}}(t,A,f,\beta_{1},\beta_{2},D)=\\ &\left\{\begin{aligned} &A&&\quantity(\displaystyle\frac{T}{2}-T_{\mathrm{a1}}<t\leq\frac{T}{2}+T_{\mathrm{a2}})&\\ &A\cos\quantity{\displaystyle\frac{2\pi f}{\beta_{1}}\quantity(\quantity|t-\frac{T}{2}|-T_{\mathrm{a1}})}&&\quantity(\displaystyle T_{\mathrm{b1}}<t\leq\frac{T}{2}-T_{\mathrm{a1}})&\\ &A\cos\quantity{\displaystyle\frac{2\pi f}{\beta_{2}}\quantity(\quantity|t-\frac{T}{2}|-T_{\mathrm{a2}})}&&\quantity(\displaystyle\frac{T}{2}+T_{\mathrm{a2}}<t\leq T-T_{\mathrm{a1}})&\\ &-A&&\text{otherwise,}&\end{aligned}\right.\end{split}$ (2)
where $f$ is the respiratory frequency, $A$ is the amplitude, $\beta_{1}$ and $\beta_{2}$ are the roll-off rates, and $D$ is the duty ratio. Note that $\beta_{1}$, $\beta_{2}$, and $D$ correspond to the temporal ratios of the expiration, inspiration, and expiratory plateau phases, respectively. In addition, $T_{\mathrm{a1}},T_{\mathrm{a2}},T_{\mathrm{b1}},$ and $T_{\mathrm{b2}}$ in (2) are determined as follows:
$T_{\mathrm{a1}}=\frac{D(1-\beta_{1})}{2f},\,T_{\mathrm{b1}}=\frac{(1-D)(1-\beta_{1})}{2f},$ (3)
$T_{\mathrm{a2}}=\frac{D(1-\beta_{2})}{2f},\,T_{\mathrm{b2}}=\frac{(1-D)(1-\beta_{2})}{2f}.$ (4)
Example MRCW models are shown in Fig. 2 for $f=1$ Hz and $A=1$. As the figure shows, the waveform for the MRCW model is dependent on the parameters $\quantity(f,\beta_{1},\beta_{2},D)$. Hereafter, the parameters $\bm{p}=\quantity(f,\beta_{1},\beta_{2},D)^{\mathrm{T}}$ are used as the respiratory features.
Figure 1: Example of respiratory displacement when measured using a radar system.
Figure 2: Examples of the respiratory displacement MRCW model with parameters (a) $(\beta_{1},\beta_{2},D)=(0.5,0.5,0.5)$, (b) $(\beta_{1},\beta_{2},D)=(0.1,0.5,0.5)$, (c) $(\beta_{1},\beta_{2},D)=(0.5,0.1,0.5)$, and (d) $(\beta_{1},\beta_{2},D)=(0.5,0.5,0.8)$.
To determine the MRCW model parameters from the measured signal, the following operation is performed:
$\begin{split}\quantity(\hat{\tau}(t),\hat{A}(t),\hat{\bm{p}}(t))&=\\ \underset{(\tau,A,\bm{p})}{\operatorname{argmin}}\int_{0}^{T_{0}}&\left|d(t^{\prime})w(t^{\prime}-t)-d_{\mathrm{MRCW}}\quantity(t^{\prime}-\tau,A,\bm{p})\right|^{2}\,{\mathrm{d}}t^{\prime},\end{split}$ (5)
$\qquad\hat{d}_{\mathrm{MRCW}}(t)$ $=d_{\mathrm{MRCW}}\quantity(t-\hat{\tau}(t),\hat{A}(t),\hat{\bm{p}}(t)),$ (6)
where $\tau$ is the time shift and $\hat{d}_{\mathrm{MRCW}}(t)$ is the MRCW model that provides the best approximation of $d(t)$. In addition, $w(t)$ is a rectangular window with a length of 8 s. The parameters estimated in (5) are defined here as the estimated waveform features $\hat{\bm{p}}(t)=(\hat{f}(t),\hat{\beta}_{1}(t),\hat{\beta}_{2}(t),\hat{D}(t))^{\mathrm{T}}$ at a time $t$.
The measured displacement waveform $d(t)$ and the estimated waveform from the MRCW model $\hat{d}_{\mathrm{MRCW}}(t)$ are shown in Fig. 3. This comparison indicates that $\hat{d}_{\mathrm{MRCW}}(t)$ approximates $d(t)$.
Figure 3: Displacement waveform measured using radar (black solid line) and calculated using the MRCW model (red dashed line).
Note that the MRCW model parameters in Fig. 3 are $\hat{f}=0.4$ Hz and $(\hat{\beta}_{1},\hat{\beta}_{2},\hat{D})=(0.9,0.6,0.9)$.
II-B Features of the Respiratory Expiratory Plateau
Although the MRCW model can approximate the respiratory displacement, the expiratory plateau waveform is modeled as a constant straight line, which is inaccurate, as illustrated in Fig. 3. In this study, the displacement waveform during the expiratory plateau time interval is used as an additional feature.
For $t$, parabolic fitting is performed using the parabolic curve $d_{\mathrm{par}}(t;c_{0},c_{1},c_{2})=c_{2}t^{2}+c_{1}t+c_{0}$ as shown:
[…]
where $w_{\mathrm{E}}^{(t)}(t^{\prime})$ is a window function that is defined as follows:
$w_{\mathrm{E}}^{(t)}(t^{\prime})=\begin{cases}w(t^{\prime}-t)&\quantity(\hat{d}_{\mathrm{MRCW}}(t^{\prime})\geq\varepsilon\hat{A}(t)),\\ 0&(\text{otherwise}),\end{cases}$ (8)
where $\varepsilon=0.6$ is a threshold that was set empirically. In this paper, $\hat{c}_{2}(t)$ is used as a feature that represents the expiratory plateau’s shape. Fig. 4 shows the time interval for the expiratory plateau (gray area) and the corresponding estimated parabolic curve (red solid line).
Figure 4: Displacement waveform measured using radar (black solid line) and calculated using the parabolic model (red solid line).
II-C Features of the Respiratory Statistics
In Section II-A, a method to estimate the respiratory features at each instant $t$ was presented. However, respiration includes both voluntary and involuntary components that result in temporal fluctuations in the respiratory feature. These temporal variations can have an adverse effect on the accuracy of individual identification. To mitigate this variability, we extract the statistical properties of the respiratory features.
First, the MRCW parameters $(\hat{f}(t),\hat{\beta}_{1}(t),\hat{\beta}_{2}(t),\hat{D}(t))$ are transformed into a six-dimensional vector $\bm{q}(t)$, as follows:
$\bm{q}(t)=\begin{bmatrix}q_{1}\\ q_{2}\\ q_{3}\\ q_{4}\\ q_{5}\\ q_{6}\end{bmatrix}=\begin{bmatrix}\hat{f}(t)\\ \hat{D}(t)\\ \hat{\beta}_{1}(t)+\hat{\beta}_{2}(t)\\ |\hat{\beta}_{1}(t)-\hat{\beta}_{2}(t)|\\ \quantity{\hat{\beta}_{1}(t)+\hat{\beta}_{2}(t)}/\hat{f}(t)\\ \hat{c}_{2}(t)\end{bmatrix}.$ (9)
In (9), $q_{1}$ and $q_{2}$ represent the respiratory frequency and the expiratory plateau duration, respectively. Furthermore, $q_{3}$ and $q_{4}$ represent the sum and the difference, respectively, of the inspiratory and expiratory rates. Additionally, $q_{5}$ represents the ratio of the inspiratory and expiratory rates to the sum of the inspiratory and expiratory rates relative to the respiratory frequency. Finally, $q_{6}$ is a coefficient of the parabolic curve estimated using (7).
Next, statistics are extracted for each element of the six-dimensional feature vector $\bm{q}(t)$ and transformed into a 24-dimensional feature vector $\bm{r}_{\mathrm{resp}}$ as $\bm{r}_{\mathrm{resp}}=[\bm{r}_{1},\bm{r}_{2},\ldots,\bm{r}_{6}]^{\mathrm{T}}\in\mathbb{R}^{24\times 1}$, where $\bm{r}_{n}=[\mu_{n},\sigma_{n},\gamma_{n},\kappa_{n}]\in\mathbb{R}^{1\times 4}$ is defined for $n=1,2,\ldots,6$; here, $\mu_{n}$, $\sigma_{n}$, $\gamma_{n}$, and $\kappa_{n}$ represent the mean, the standard deviation, the skewness, and the kurtosis of $q_{n}$, respectively.
III Heartbeat features for individual identification
III-A High-Frequency Enhancement Using Differentiation
In this section, we propose a method to extract the features of the heartbeat component required for individual identification. Because the fundamental frequency component of the heartbeat is often masked by harmonic components of the respiration, it has proven to be advantageous to use the higher harmonics of the heartbeat components [18]. Therefore, our proposed method emphasizes the high-frequency components by using second-order differentiation, as in our previous work [21]. Because the second-order differentiation $s^{\prime\prime}(t)=({\mathrm{d}}^{2}/{\mathrm{d}}t^{2})s(t)$ is expressed as $-\omega^{2}S(\omega)$ in the frequency domain, the high-frequency components, including the harmonics of the heartbeat, are emphasized; here, $\omega$ is the angular frequency and $S(\omega)$ represents the Fourier transform of $s(t)$.
III-B Feature Extraction Using Mel-Frequency Cepstrum
Next, we use the mel-frequency cepstral coefficients (MFCCs) to generate features. Historically, MFCCs have been used to convert the short-term spectra of speech signals into feature vectors [53]. Because radar-measured heartbeat signals are periodic within a short time, in exactly the same manner as speech signals, use of MFCCs to generate heartbeat features is appropriate when using radar [47]. In [47], MFCCs were applied to real signals that were estimated from the phases of reflected signals. However, our previous work [21] has shown that signal processing using complex reflected signals directly improves the heartbeat estimation accuracy. Therefore, in this section, we propose a method to extract heartbeat features by applying MFCCs to complex signals. The calculation of the MFCCs is performed using the following steps.
III-B1 Short-time Fourier transform
For the differentiated signal $s^{\prime\prime}(t)$, the short-time Fourier transform (STFT) is applied as follows:
$S(t,f)=\int_{-\infty}^{\infty}s^{\prime\prime}(\tau)w_{\mathrm{STFT}}(\tau-t)\exp\left(-\,\mathrm{j}2\pi f\tau\right){\mathrm{d}}\tau,$ (10)
where $w_{\mathrm{STFT}}(t)$ is a rectangular window function with a width of $T_{\mathrm{STFT}}$. This window width $T_{\mathrm{STFT}}$ should be greater than the heartbeat interval and we set $T_{\mathrm{STFT}}=2.0$ s empirically.
III-B2 Mel filters
Next, we apply a mel filter bank to $S(t,f)$ to calculate the mel-frequency spectrum. The $\ell$th mel filter $H_{\ell}(f)$ is expressed as:
$H_{\ell}(f)$ =$=\begin{cases}\vspace{2mm}\displaystyle\frac{2(f-f_{\ell})}{(f_{\ell+1}-f_{\ell})(f_{\ell+2}-f_{\ell})}&\displaystyle(f_{\ell}\leq f<f_{\ell+1}),\\ \vspace{2mm}\displaystyle\frac{2(f_{\ell+2}-f)}{(f_{\ell+2}-f_{\ell+1})(f_{\ell+2}-f_{\ell})}&\displaystyle(f_{\ell+1}\leq f<f_{\ell+2}),\\ 0&\text{otherwise}\end{cases}$ (11)
for $\ell=0,1,\ldots,L-1$, where
$f_{\ell}$ $=\tilde{f}\quantity{\exp\quantity(\frac{m_{\ell}}{\tilde{m}})-1},$ (12)
is the center frequency of the $\ell$th mel filter corresponding to the $\ell$th mel scale. Here, the mel scale $m_{\ell}$ is defined as
$m_{\ell}=\tilde{m}\frac{\ell}{L+1}\log\left(1+f_{\mathrm{s}}/2\tilde{f}\right).$ (13)
for $\ell=0,1,\ldots,L+1$, where $f_{\mathrm{s}}$ is the sampling frequency and $\tilde{m}$ is defined as:
$\tilde{m}$ $=\frac{f^{\prime}}{\log\quantity(\displaystyle f^{\prime}/\tilde{f}+1)}.$ (14)
An example of a mel filter bank and the power spectrum of $s^{\prime\prime}(t)$ are shown in Fig. 5 for the case where $\tilde{f}=5.0$ Hz, $f^{\prime}=1.0$ kHz, and $L=64$, where these parameters were set following the convention in [47].
Figure 5: Examples of the mel filters and the power spectrum for $s^{\prime\prime}(t)$.
In this study, we apply the mel filters to a complex radar signal, where the mel filters are applied to the spectrum for both positive and negative frequencies as follows:
$S_{\ell}$ $=\int_{0}^{T_{0}}\int_{0}^{f_{\mathrm{s}}/2}|S(t,f)|H_{\ell}(f)\,{\mathrm{d}}f{\mathrm{d}}t,$ (15)
$S_{-\ell}$ $=\int_{0}^{T_{0}}\int_{-f_{\mathrm{s}}/2}^{0}|S(t,f)|H_{\ell}(-f)\,{\mathrm{d}}f{\mathrm{d}}t$ (16)
for $10\leq\ell\leq L-$ 1, where $T_{0}$ is the measurement time. Note that in this case, $S_{+0}$ and $S_{-0}$ are regarded as different variables.
III-B3 Discrete cosine transform
Next, we obtain the MFCC $C_{k}$ by applying a discrete cosine transform (DCT) to $\log S_{\ell}$ as shown here:
$C_{k}=$ $\frac{2}{L+\delta_{k0}}\sum_{\ell=0}^{L-1}\log S_{\ell}\cos\quantity{\frac{(2\ell+1)k\pi}{K}}$ (17)
for $k=0,1,\ldots,K-1$, and
$C_{k}=$ $\frac{2}{L+\delta_{k0}}\sum_{\ell=0}^{L-1}\log S_{-\ell}\cos\quantity{\frac{(2\ell+1)k\pi}{K}}$ (18)
for $k=-0,-1,-2,\ldots,-(K-1)$, where $\delta_{i,j}$ is the Kronecker delta, $K$ is the cepstrum dimension, and we set $K=64$. Note that $C_{+0}$ and $C_{-0}$ are also regarded as different variables here.
Finally, from the $2K$-dimensional MFCC $C_{k}$, the lower order coefficients up to $2K^{\prime}$ dimensions are extracted, where the condition $0<K^{\prime}<K$ holds; the feature vector of the heartbeat $\bm{r}_{\mathrm{hb}}=[C_{-(K^{\prime}-1)},C_{-(K^{\prime}-2)},\ldots,C_{-0},C_{+0},\ldots,C_{K^{\prime}-1}]^{\mathrm{T}}$ is defined and is used to perform individual identification in the following section, where we set $K^{\prime}=24$ empirically.
IV Performance Evaluation of the Proposed Method
In this section, we evaluate the individual identification performance of the proposed method via comparison with conventional methods. To identify an effective set of features, we compare the performance metrics obtained when the proposed feature vectors are used separately and together. We compare three cases: (1) use of the respiratory feature vector $\bm{r}_{\mathrm{resp}}$ only; (2) use of the heartbeat feature vector $\bm{r}_{\mathrm{hb}}$ only; and (3) use of both vectors as $\bm{r}_{\mathrm{prop}}=[\bm{r}_{\mathrm{resp}}^{\mathrm{T}},\bm{r}_{\mathrm{hb}}^{\mathrm{T}}]^{\mathrm{T}}$. We also compare three classifiers: the support vector machine (SVM), $k$-nearest neighbors ($k$-NN), and multilayer perceptron (MLP) classifiers. For these three classifiers, we use the three feature vector types, i.e., $\bm{r}_{\mathrm{resp}}$, $\bm{r}_{\mathrm{hb}}$, and $\bm{r}_{\mathrm{prop}}$, to compare a total of nine different approaches. Table I summarizes the definitions of nine methods for comparison; for example, method A1 indicates the identification process when using $\bm{r}_{\mathrm{resp}}$ and the SVM.
TABLE I: DEFINITIONS OF METHODS FOR INDIVIDUAL IDENTIFICATION
Method Signal Vital sign Feature Classifier
Method A1 $\angle s(t)$ Respiration $\bm{r}_{\mathrm{resp}}$ SVM
Method A2 $k$-NN
Method A3 MLP
Method B1 $\frac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}s(t)$ Heartbeat $\bm{r}_{\mathrm{hb}}$ SVM
Method B2 $k$-NN
Method B3 MLP
Method C1 $\angle s(t)$ $\frac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}s(t)$ Respiration & heartbeat $\bm{r}_{\mathrm{prop}}$ SVM
Method C2 $k$-NN
Method C3 MLP
IV-A Experiment 1: Five-Day Measurement of Six Participants
In this experiment, measurements were conducted over five days with six healthy adult participants. Each day, measurements were taken five times per participant in the morning and another five times per participant in the afternoon, with each session lasting for $60.0~{}\mathrm{s}$. Therefore, the total number of measurements performed was $(5+5)\times 5=50$ times per each person, giving a total of 300 measurements. All participants were in good health with no history of respiratory problems or heart disease. All measurements were conducted with each participant seated in front of the radar system at a distance of $1.5~{}\mathrm{m}$. The experimental setup is pictured in Fig. 6.
Figure 6: Experimental setup with the radar system and a seated participant.
The radar specifications are listed in Table II. We used a frequency-modulated continuous-wave (FMCW) radar system with a center frequency of $79.0~{}\mathrm{GHz}$ and a bandwidth of $3.6~{}\mathrm{GHz}$. The antenna array consists of three transmitting and four receiving elements, which form a MIMO array that can be approximated virtually as a 12-element linear array. The range and the direction of arrival of the echo from the target are identified by performing a Fourier transform in terms of the fast time  [54] and the virtual element number corresponding to the beamformer method [55].
TABLE II: SPECIFICATIONS OF THE RADAR SYSTEM
Parameter Value
Modulation FMCW
Center frequency $79.0~{}\mathrm{GHz}$
Center wavelength $3.8~{}\mathrm{mm}$
Bandwidth $3.6~{}\mathrm{GHz}$
No. of transmitting (Tx) antennas $3$
No. of receiving (Rx) antennas $4$
Tx element spacing $7.6~{}\mathrm{mm}$
Rx element spacing $1.9~{}\mathrm{mm}$
Beamwidths of Tx elements (E-/H-planes) ±$\pm 4^{\circ}/\pm 33^{\circ}$
Beamwidths of Rx elements (E-/H-planes) ±$\pm 4^{\circ}/\pm 45^{\circ}$
Range resolution $44.7~{}\mathrm{mm}$
Sampling frequency (slow-time) $100~{}\mathrm{Hz}$
The feature vectors $\bm{r}_{\mathrm{resp}}$, $\bm{r}_{\mathrm{hb}}$, and $\bm{r}_{\mathrm{prop}}$ are visualized by compressing them into two dimensions using the t-distributed stochastic neighbor embedding (t-SNE) method, [56] as shown in Fig. 7, which shows that the feature vectors form clusters corresponding to the six participants. In particular, when compared with the other feature vectors, the clusters are separated clearly when using $\bm{r}_{\mathrm{prop}}$. In the $\bm{r}_{\mathrm{resp}}$ case, the cluster for participant 6 is erroneously split into two separate clusters and also overlaps with other clusters.
Figure 7: Two-dimensional t-SNE visualizations of the feature vectors (a) $\bm{r}_{\mathrm{resp}}$, (b) $\bm{r}_{\mathrm{hb}}$, and (c) $\bm{r}_{\mathrm{prop}}$.
The accuracy during individual identification was evaluated using a stratified $k$-fold cross-validation approach with $k=10$, thus ensuring that each subset maintained the class distribution of the original dataset; each cross-validation iteration used 270 samples for training and 30 samples for validation. As a single-fold validation, only data acquired from the same period can be used as validation data. For example, in the single-fold method, data measured on the morning of the first day are used as the validation data and the remaining data serve as training data.
The three feature vectors $\bm{r}_{\mathrm{resp}}$, $\bm{r}_{\mathrm{hb}}$, and $\bm{r}_{\mathrm{prop}}$ are input into the three classifiers to compare their identification accuracies. Here, because both SVM and $k$-NN are binary classifiers, the one-versus-rest method is used to perform multi-class classification [57]. In addition, the activation function used for the MLP output layer is the softmax function. The parameters for each classifier are optimized via a grid search. The optimization processes are performed for each of the three feature vectors $\bm{r}_{\mathrm{resp}}$, $\bm{r}_{\mathrm{hb}}$, and $\bm{r}_{\mathrm{prop}}$ individually. The parameters to be optimized and the values selected are given in Table III.
TABLE III: OPTIMIZED PARAMETERS FOR EACH METHOD
   Classifier    Parameters Options Feature vector
A) $\bm{r}_{\mathrm{resp}}$ B) $\bm{r}_{\mathrm{hb}}$ C) $\bm{r}_{\mathrm{prop}}$
1) SVM Kernel function Linear / Gaussian Gaussian Gaussian Gaussian
2) $k$-NN No. of neighbors $[1,300]$ 9 28 28
Distance Euclidean, Cityblock, Cosine Cityblock Cosine Cosine
3) MLP No. of layers $[1,3]$ 2 2 2
[…]
Activation function ReLU, Sigmoid ReLU ReLU ReLU
  • : Rectified linear unit
The confusion matrices obtained for each classifier are shown in Fig. 8. The accuracy indicates the proportion of correctly identified cases with respect to the total number of cases. When using $\bm{r}_{\mathrm{resp}}$, $\bm{r}_{\mathrm{hb}}$, and $\bm{r}_{\mathrm{prop}}$, these three classifiers achieved accuracies of approximately 80%, 90%, and 95%, respectively. These results suggest that the heartbeat feature is superior to the respiratory feature when used in individual identification, and the combination of the respiratory and heartbeat components is superior to both the respiratory feature and the heartbeat feature when they are used separately. In Fig. 8, we see that method C1 realizes the highest accuracy of 96.33% from among the nine methods.
Figure 8: Confusion matrices for each method. (a) SVM, (b) $k$-NN, (c) MLP.
Fig. 9 presents the receiver operating characteristic (ROC) curves and the corresponding area under the curve (AUC) results for each classifier. Note that the ROC curves shown in Fig. 9 are the macro-averages of the one-versus-rest ROC curves for each class. Fig. 9 shows that when $\bm{r}_{\mathrm{prop}}$ is used, the AUC for all classifiers reaches 0.99 and is higher than that realized when using $\bm{r}_{\mathrm{resp}}$ and $\bm{r}_{\mathrm{hb}}$ separately.
Table IV presents the F1 scores and the AUCs for each participant when $\bm{r}_{\mathrm{prop}}$ is used as the input, along with their respective macro-averages. From Table IV, we see that the SVM achieves the highest F1 scores and AUCs among the classifiers. The AUC exceeds a macro average of 0.99 for all classifiers, thus indicating that each classifier achieves comparable accuracy to the others. This indicates that the proposed features enable high-accuracy identification, regardless of the classifier that is used.
Figure 9: ROC curves for (a) SVM, (b) $k$-NN, and (c) MLP.
TABLE IV: F1 SCORE AND AUC FOR EACH PARTICIPANT USING $\bm{r}_{\mathrm{prop}}$ IN EXP. 1
  Classifier   Participant   F1 score   AUC
1) SVM (method C1) 1 0.959 0.994
2 0.990 0.999
3 0.904 0.995
4 0.969 1.000
5 0.961 0.999
6 1.000 1.000
macro-ave. 0.964 0.998
2) $k$-NN (method C2) 1 0.938 0.998
2 0.960 0.997
3 0.898 0.996
4 0.961 0.998
5 0.952 0.999
6 0.990 1.000
macro-ave. 0.950 0.998
3) MLP (method C3) 1 0.951 0.999
2 0.960 0.997
3 0.860 0.979
4 0.970 0.999
5 0.923 0.997
6 0.990 0.999
macro-ave. 0.942 0.995
IV-B Experiment 2: Public Dataset of Thirty Participants
Next, we evaluated the accuracy of the proposed method using a public dataset provided by Schellenberger et al. [58], where they used a continuous-wave (CW) radar system with an operating frequency of $24~{}\mathrm{GHz}$ to collect echoes from a total of 30 individuals, comprising 14 male and 16 female participants. The echoes were recorded with a sampling frequency of $2.0\,\mathrm{kHz}$, which was then downsampled to $100\,\mathrm{Hz}$ to make the conditions consistent with those of experiment 1. The measurement time varied for each participant and ranged from approximately $600\,\mathrm{s}$ to $700\,\mathrm{s}$. In this study, we divided the data into segments with a length of $T_{0}$ and examined two cases, where $T_{0}=60.0\,\mathrm{s}$ and $T_{0}=5.0\,\mathrm{s}$; a feature vector was then generated from each segment,
Fig. 10 visualizes two-dimensional plots of the feature vector $\bm{r}_{\mathrm{hb}}$ as obtained from data segments with $T_{0}=60.0\,\mathrm{s}$ and $T_{0}=5.0\,\mathrm{s}$ when compressed using t-SNE. In both cases, it was observed that almost all plots were clearly separated into distinct clusters corresponding to the 30 participants. Next, the feature vector $\bm{r}_{\mathrm{hb}}$ obtained was input into the SVM and the identification accuracy was evaluated by performing a stratified $k$-fold cross-validation with $k=5$; the resulting confusion matrices are shown in Fig. 11. First, when $T_{0}=60.0\;\mathrm{s}$, 100% accuracy was achieved for all but four participants, and the overall accuracy was 98.69%, thus indicating that individual identification was achieved successfully. Next, when $T_{0}=5.0\,\mathrm{s}$, the overall accuracy realized was 99.39%, which also represents high-precision identification.
Figure 10: Two-dimensional plots of feature vectors $\bm{r}_{\mathrm{hb}}$ compressed by t-SNE when using the public dataset with a data length (a) $T_{0}=60.0$ s and (b) $T_{0}=5.0$ s.
Figure 11: Confusion matrices for experiment 2 with a data length (a) $T_{0}=60\;\mathrm{s}$ and (b) $T_{0}=5\;\mathrm{s}$.
In experiment 1, the accuracy of method B2 was only 93.33%, whereas in experiment 2, the same method B2 achieved accuracy of 98.69%, partly because the public dataset was collected within a day and was thus unaffected by daily variations in the physiological signals. We should also note that the accuracy obtained for $T_{0}=5.0\;\mathrm{s}$ was almost the same as the accuracy for $T_{0}=60.0\;\mathrm{s}$, which may indicate that long measurements are not necessary during individual identification if the daily variations in the physiological signals can be neglected.
V Discussion
Here we compare the results from this study with those from existing publications, as shown in Table V. Although these studies assumed different measurement conditions (e.g., sensor types, the number of antennas used, transmitting power, distance between the antenna and the target person), we can see that the proposed method achieves an accuracy that is comparable to that reported in the existing studies.
In particular, the work of Peng et al. [47] is similar to that of this study because they also used the respiratory and heartbeat features for individual identification; they used the respiratory features for individual identification, and they also used the heartbeat features only during sleep apnea events because the respiratory signals cannot be obtained during apnea events. In contrast, this study improves the overall identification accuracy by using a combination of the respiratory and heartbeat features simultaneously, which differentiates this study from that of Peng et al. [47].
TABLE V: COMPARISON WITH RELATED WORKS
Study Vital sign System No. of Accuracy
subjects
Hwang et al. [46] HB CW 30 88.9%
Peng et al. [47] Resp. & HB FMCW 5 94.4%
Korany et al. [28] Resp. Wi-Fi 20 95.0%
Wang et al. [42] Resp. FMCW 37 96.0%
Lin et al. [49] HB CW 78 98.6%
This study Resp. & HB FMCW 6 96.3%
HB CW 30 99.4%
VI Conclusion
In this study, we developed a radar-based individual identification method using both respiratory and heartbeat features. The proposed method extracts the subject’s respiratory features using a modified raised-cosine-waveform model. In addition, a parabolic model is used to express the waveform during the expiratory plateau phase. The proposed method extracts the heartbeat features by applying a time-frequency analysis and a mel-frequency cepstral analysis to the second time-derivative of the complex radar signal. To identify a suitable combination of features and a classifier, we compared the performances of nine methods based on combinations of three feature vectors, comprising the respiratory feature, the heartbeat feature, and the respiratory and heartbeat feature, and three classifiers, comprising the support vector machine, $k$-nearest neighbors, and multilayer perceptron methods. The accuracy of the proposed method during individual identification was evaluated in two scenarios; one scenario involved five-day radar measurements of six participants conducted by the authors, and the other scenario involved one-day radar measurement of thirty participants, which was provided as a public dataset. Through performance evaluation, we demonstrated that the combination of the respiratory and heartbeat features achieved accuracy of 96.33% for the five-day radar data from six participants, and accuracy of 99.39% was realized for a public dataset composed of thirty participants using a short data segment that lasted for only 5.0 s; these results indicate the effectiveness of the proposed approach for radar-based individual identification applications.
ACKNOWLEDGMENT
This work was supported in part by SECOM Science and Technology Foundation, in part by JST under Grants JPMJMI22J2 and JPMJMS2296, in part by JSPS KAKENHI under Grants 21H03427, 23H01420, 23K19119, and 24K17286, and in part by the New Energy and Industrial Technology Development Organization (NEDO). This work involved human subjects in its research. Approval of all ethical and experimental procedures and protocols was granted by the Ethics Committee of the Graduate School of Engineering, Kyoto University, under Approval No. 202223.